OpenStudy (anonymous):

Use the endpoint "R" and midpoint "M" or RS (line segment) to find the coordinates of the other endpoint "S" R(-3,-2) , M(-1,-8)

OpenStudy (anonymous):

This is geometry

OpenStudy (anonymous):

I would appreciate if you show the work It would help me with the other problems. Thanks

OpenStudy (cwtan):

Let S(x,y) Midpoint =\(\Huge\frac{x_1+x_2}{2},\frac {y_1+y_2}{2}\) (x-3)/2=-1 (y-2)/2=-8

OpenStudy (anonymous):

thanks :)